# Virtual Reality User Interface Design: Best Practices and Implementation

### v1

@@ 3.2.2. VR Game Design & Development
Çatak et al. (Çatak et al., 2020) provide comprehensive guidelines for VR game design, covering perceptual design principles such as visual attraction points, as well as depth and sensory cue utilization. Their work also addresses interactional and navigational design guidelines. Dombrowski et al. (Dombrowski et al., 2019) focus on inclusive VR game design through the pillars of universal design, aiming to ensure equitable use for users with diverse abilities and preferences, as well as immersive and accessible content that minimizes physical effort.
Anuyahong (Anuyahong and Pengnate, 2023) demonstrates that there is a strong relationship between effective UI design and positive user experiences while emphasizing the importance of intuitive navigation and consistent visual design. Wang et al. (Wang et al., 2021b) examine authenticity, interactivity, and collaboration in VR games, providing best practices for creating engaging experiences that support learning and collaborative problem-solving in games. Ap Cenydd and Headleand (ap Cenydd and Headleand, 2018) present movement modalities in VR, offering guidelines for player guidance, circular environment design, and comfort preservation through appropriate character distancing and exploration rewards.
Zhang et al. (Zhang et al., 2025) contribute specialized guidelines for inclusive avatar design that support disability representation in social VR environments, addressing avatar body appearance, dynamics, assistive technology integration, and interface considerations for creating accessible and inclusive virtual experiences. Puspitasari and Lee (Puspitasari and Lee, 2022) review persuasive user interface strategies for addressing addiction in virtual environments, providing guidelines for implementing break encouragement systems, user statistics tracking, and adaptive aesthetics to promote responsible VR use.
3.2.3. Onboarding & Learning
@@
Chauvergne et al. (Chauvergne et al., 2023) investigate practices in VR user onboarding, emphasizing the importance of instruction availability, corrective feedback, and context-based learning approaches that integrate instructions into meaningful tasks. Their research highlights the need to allocate time for familiarizing users with the controls before interacting with the virtual environment. Gonzalez-Argote and Gonzalez-Argote (Gonzalez-Argote and Gonzalez-Argote, 2023) establish best practices for immersive learning design, focusing on aligning immersive learning experiences with clear learning objectives, selecting appropriate technology tailored to the target audience, and designing interactive, engaging environments that support meaningful exploration, collaboration, realistic simulations, and progress tracking through feedback and assessment.
3.2.4. Menus & Interface Elements
@@
Azai et al. (Azai et al., 2018a) introduce the Open Palm Menu concept, providing guidelines for positioning menu items close to users’ hands and designing vertical menu layouts for optimal selection efficiency. Li et al. (Li et al., 2024) offer recommendations for customizable menu configurations and body-area-specific icon placement based on application type. Lou et al. (Lou et al., 2021) present hand-adaptive user interface best practices that emphasize placing interactive objects at lower heights to reduce fatigue and leveraging spatial memory for intuitive object positioning. Lediaeva and Jr (Lediaeva and LaViola, 2020) provide specific recommendations for arm-mounted, waist-level, hand-based, and spatial menus with corresponding selection techniques.
Wentzel et al. (Wentzel et al., 2024) compare virtual reality menu archetypes, offering design recommendations across different menu layouts, item counts, and levels of hierarchy. Wang et al. (Wang et al., 2021a) investigate menu selection methods in immersive virtual environments, comparing fixed and handheld menu approaches, and provide guidelines for choosing appropriate selection techniques. Dingler et al. (Dingler et al., 2018) assess text parameters for reading in VR, providing specific guidelines for angular height, vergence distance, viewbox parameters, and font selection to optimize text readability and user comfort in virtual environments.
3.2.5. Cybersickness & User Comfort
@@
Ramaseri Chandra et al. (Ramaseri Chandra et al., 2022) provide design guidelines that address latency requirements, movement realism, visual flicker prevention, logical environment design, optimal field of view, and session duration. Yahaya et al. (Yahaya et al., 2022) present cybersickness reduction guidelines, offering specific recommendations for duration management, user experience preparation, physical setting optimization, movement control strategies and more.
3.3. Integration
@@
3.4. Implementation
@@
Figure 2. Tutorial highlights: (a) UI placed 3 meters away for the user. (b) UI with no floating elements. (c) Rounded visual design. (d) Buttons placed at lower heights.
[…]
The figure shows four parts of the tutorial: distance, absence of floating elements, rounded design, and button placement.
Figure 3. Design one: A settings menu built according to our VR UI guidelines.
[…]
The figure shows a VR user settings menu designed with the best practices in mind.
3.4.1. Design One
A user settings menu built according to the best practices. Users can interact with the interface by controller ray-cast. Selection happens by pressing the front trigger button of the controller and the joystick is used for scrolling. As shown in Figure 3, key features of the interface include a comfortable viewing distance, content centered within the user’s field of view, and interactive elements such as buttons and toggles placed at lower heights for ergonomic access. The interface provides visual, audio, and haptic feedback, uses vertical placement for interface elements, and includes instructions that remain accessible at all times. It maintains a consistent color scheme, features easily readable text, and avoids blur, floating elements, irregular shapes, or unbalanced backgrounds.
3.4.2. Design Two
A user settings menu designed to intentionally violate several of the proposed VR UI design guidelines. Interaction is handled through hand-tracking ray-cast, with selection performed by pinching the thumb and index finger, and scrolling achieved through a pinch-and-drag gesture. As demonstrated in Figure 4, in this interface, elements are misaligned, placed too close to the user, and positioned in their peripheral vision. Feedback is minimal or inconsistent, with no haptic feedback provided.
@@ 3.4.2. Design Two
Figure 4. Design Two: A deliberately poorly designed settings menu that does not adhere to the guidelines.
[…]
The figure shows a poorly designed VR settings menu that does not adhere to the guidelines.
3.4.3. Tutorial
Table 1. Survey questions used to evaluate each interface
@@ 3.5. User Study
The survey consisted of 12 Likert-scale questions rated from 1 to 5 (worst to best) assessing aspects such as usability, comfort, clarity, ease of interaction, and overall experience. In addition, participants were provided with an open-ended field where they could freely write any observations, thoughts, or preferences regarding each design. After completing both interactions and surveys, participants were asked to choose which design they preferred overall.
The full list of survey questions is provided in Table 1.
4. Results
@@
4.1. Quantitative Analysis
Design One consistently received higher ratings across all questions in both exposure orders. Figure 5 presents the mean ratings for each survey question, comparing the performance of Design One and Design Two. The values represent the combined average scores for each design, based on responses from both Group A and Group B, regardless of whether the design was experienced first or second.
Figure 5. Mean Likert scores (1–5) for each of the 12 evaluation questions, comparing Design One and Design Two.
[…]
The figure shows mean scores for each survey question, demonstrating higher scores for Design One for all 12 questions.
@@ 4.1. Quantitative Analysis
When experienced first with Group A, Design One averaged a rating of 4.6 out of 5 across all questions. With Group B, it maintained a similar high performance with a slightly lower average rating of 4.5, indicating a consistently positive user experience, even after initial exposure to a less optimal interface. Figure 6(a) shows a slight decline in Q6 (Feedback), Q7 (Visual Appeal), and Q9 (Color Comfort). This can be attributed to some users either not noticing the haptic feedback or finding it unremarkable, as well as a few comments suggesting the design felt visually plain or lacking color. Q11 (Text Readability) shows a noticeable increase when Design One was experienced second, indicating that users in Group B found it easier to read, especially in contrast to the less readable text in Design Two.
(a) Design One
[…]
The figure shows mean scores for each survey question for Design One for both the first and second interaction order. A few questions have slightly lower scores for the second interaction compared to the first.
(b) Design Two
[…]
The figure shows mean scores for each survey question for Design Two for both the first and second interaction order, demonstrating higher scores for the first interaction compared to lower scores when Design Two was tested second.
Figure 6. Mean survey scores (1-5) by question showing that Design One stays consistently high regardless of whether users saw it first or second, whereas Design Two’s scores fall significantly, particularly for comfort (Q2), interaction ease (Q3), and element organization (Q4), when experienced after the superior design.
In contrast, Design Two performed worse overall, with noticeable effects based on the order of exposure. When presented first, it received an average score of 3.4 across all questions. However, when experienced after Design One, the average rating of Group A dropped significantly to 2.9, indicating that prior exposure to a better-designed interface made users more aware of the flaws in the second design. As shown in Figure 6(b), the most significant drops were observed in Q2 (Comfort), Q3 (Ease of Interaction), and Q4 (Organization of Elements), suggesting that users found Design Two notably less comfortable, harder to use, and poorly structured when compared directly to a well-designed interface.
4.2. Overall Preference

### v2

@@ 2.2.2 VR Game Design & Development
Çatak et al. [ccatak2020guideline] provide comprehensive guidelines for VR game design, covering perceptual design principles such as visual attraction points, as well as depth and sensory cue utilization. Their work also addresses interactional and navigational design guidelines. Dombrowski et al. [dombrowski2019designing] focus on inclusive VR game design through the pillars of universal design, aiming to ensure equitable use for users with diverse abilities and preferences, as well as immersive and accessible content that minimizes physical effort.
Anuyahong [Anuyahong_Pengnate_2023] demonstrates that there is a strong relationship between effective UI design and positive user experiences while emphasizing the importance of intuitive navigation and consistent visual design. Wang et al. [wang2021authenticity] examine authenticity, interactivity, and collaboration in VR games, providing best practices for creating engaging experiences that support learning and collaborative problem-solving in games. Ap Cenydd and Headleand [ap2018movement] present movement modalities in VR, offering guidelines for player guidance, circular environment design, and comfort preservation through appropriate character distancing and exploration rewards.
Zhang et al. [zhang2025inclusive] contribute specialized guidelines for inclusive avatar design that support disability representation in social VR environments, addressing avatar body appearance, dynamics, assistive technology integration, and interface considerations for creating accessible and inclusive virtual experiences. Puspitasari and Lee [puspitasari2022review] review persuasive user interface strategies for addressing addiction in virtual environments, providing guidelines for implementing break encouragement systems, user statistics tracking, and adaptive aesthetics to promote responsible VR use.
2.2.3 Onboarding & Learning
@@
Chauvergne et al. [chauvergne2023user] investigate practices in VR user onboarding, emphasizing the importance of instruction availability, corrective feedback, and context-based learning approaches that integrate instructions into meaningful tasks. Their research highlights the need to allocate time for familiarizing users with the controls before interacting with the virtual environment. Gonzalez-Argote and Gonzalez-Argote [gonzalez202310] establish best practices for immersive learning design, focusing on aligning immersive learning experiences with clear learning objectives, selecting appropriate technology tailored to the target audience, and designing interactive, engaging environments that support meaningful exploration, collaboration, realistic simulations, and progress tracking through feedback and assessment.
2.2.4 Menus & Interface Elements
@@
Azai et al. [azai2018open] introduce the Open Palm Menu concept, providing guidelines for positioning menu items close to users’ hands and designing vertical menu layouts for optimal selection efficiency. Li et al. [li2024investigating] offer recommendations for customizable menu configurations and body-area-specific icon placement based on application type. Lou et al. [lou2021hand] present hand-adaptive user interface best practices that emphasize placing interactive objects at lower heights to reduce fatigue and leveraging spatial memory for intuitive object positioning. Lediaeva and Jr [lediaeva2020evaluation] provide specific recommendations for arm-mounted, waist-level, hand-based, and spatial menus with corresponding selection techniques.
Wentzel et al. [wentzel2024comparison] compare virtual reality menu archetypes, offering design recommendations across different menu layouts, item counts, and levels of hierarchy. Wang et al. [wang2021experimental] investigate menu selection methods in immersive virtual environments, comparing fixed and handheld menu approaches, and provide guidelines for choosing appropriate selection techniques. Dingler et al. [dingler2018vr] assess text parameters for reading in VR, providing specific guidelines for angular height, vergence distance, viewbox parameters, and font selection to optimize text readability and user comfort in virtual environments.
2.2.5 Cybersickness & User Comfort
@@
Ramaseri Chandra et al. [ramaseri2022systematic] provide design guidelines that address latency requirements, movement realism, visual flicker prevention, logical environment design, optimal field of view, and session duration. Yahaya et al. [yahaya2022comparative] present cybersickness reduction guidelines, offering specific recommendations for duration management, user experience preparation, physical setting optimization, movement control strategies and more.
2.3 Integration
@@
  • Guideline 16.4.1: Provide suitable icons, logos, and slogans that reflect disability communities. Users should be able to add well-known icons and slogans to their avatars that represent different disability groups and affirm their identity.
@@
2.4 VR UI Guidelines Web Tool
To support practical adoption of the synthesized VR UI design guidelines, we developed a web application [^3] Web application: https://vr-guidelines.se.cit.tum.de/ that serves as an interactive guideline explorer and planning tool. While the complete guideline set is available on the project’s GitHub Wiki, the web tool provides a structured workflow for discovering, selecting, and documenting guidelines that are most relevant to a specific VR project.
The tool is implemented using HTML, CSS, and JavaScript. The guideline set can be explored through an interactive filtering interface, as shown in Figure 2.4. Users specify the application context and may additionally select application features, design aspects, design goals and priorities, and design challenges and concerns. Based on these selections, the tool returns a tailored subset of guidelines by matching the chosen categories against each guideline’s assigned tags. This allows developers and designers to quickly narrow the guidelines and focus on recommendations aligned with their intended use case (e.g., education or gaming) and design priorities (e.g., immersion or accessibility). For transparency and exploration, the tool also provides access to the complete list of guidelines without filtering.
Figure 2: Filtering interface of the VR UI Guidelines tool. Users select application context and optional categories to obtain a tailored guideline subset.
To facilitate systematic implementation in the developement process, the web tool includes a checklist where users can mark guidelines as completed as they are integrated into an application (Figure 2.4).
Figure 3: Checklist for tracking guideline implementation.
Building on this, the tool generates an implementation report summarizing overall completion percentage, counts of completed and remaining guidelines, and gives a category-level progress breakdown. This is shown in Figure 2.4. Both the checklist and the report can be exported as PDF documents, enabling teams to use the tool as documentation for design decisions, implementation progress, and internal reviews.
Overall, the website extends the contribution of this work from a static set of recommendations to an actionable, interactive tool that supports learning, project-specific guideline selection, and traceable implementation of best practices in VR UI design.
Figure 4: Automatically generated progress report summarizing overall and category-level completion.
2.5 Implementation
@@
2.5.1 Design One
A user settings menu built according to the best practices. Users can interact with the interface by controller ray-cast. Selection happens by pressing the front trigger button of the controller and the joystick is used for scrolling. As shown in Figure 2.5.1, key features of the interface include a comfortable viewing distance, content centered within the user’s field of view, and interactive elements such as buttons and toggles placed at lower heights for ergonomic access. The interface provides visual, audio, and haptic feedback, uses vertical placement for interface elements, and includes instructions that remain accessible at all times. It maintains a consistent color scheme, features easily readable text, and avoids blur, floating elements, or irregular shapes.
Figure 5: Tutorial highlights showing different UI design choices: (a) distance placement, (b) no floating elements, (c) rounded visuals, and (d) lower button placement.
Figure 6: Design one: A settings menu built according to our VR UI guidelines.
2.5.2 Design Two
A user settings menu designed to intentionally violate several of the proposed VR UI design guidelines. Interaction is handled through hand-tracking ray-cast, with selection performed by pinching the thumb and index finger, and scrolling achieved through a pinch-and-drag gesture. As demonstrated in Figure 2.5.2, in this interface, elements are misaligned, placed too close to the user, and positioned in their peripheral vision. Feedback is minimal or inconsistent, with no haptic feedback provided.
@@ 2.5.2 Design Two
Figure 7: Design Two: A deliberately poorly designed settings menu that does not adhere to the guidelines.
Table 1: Survey questions used to evaluate each interface
@@ 2.6 User Study
Figure 8: Mean Likert scores (1–5) for each of the 12 evaluation questions, comparing Design One and Design Two.
@@
The survey consisted of 12 Likert-scale questions rated from 1 to 5 (worst to best) assessing aspects such as usability, comfort, clarity, ease of interaction, and overall experience. In addition, participants were provided with an open-ended field where they could freely write any observations, thoughts, or preferences regarding each design. After completing both interactions and surveys, participants were asked to choose which design they preferred overall. The full list of survey questions is provided in Table 1.
3 Results
@@
3.1 Quantitative Analysis
Design One consistently received higher ratings across all questions in both exposure orders. Figure 2.6 presents the mean ratings for each survey question, comparing the performance of Design One and Design Two. The values represent the combined average scores for each design, based on responses from both Group A and Group B, regardless of whether the design was experienced first or second.
@@ 3.1 Quantitative Analysis
When experienced first with Group A, Design One averaged a rating of 4.6 out of 5 across all questions. With Group B, it maintained a similar high performance with a slightly lower average rating of 4.5, indicating a consistently positive user experience, even after initial exposure to a less optimal interface. Figure 3.1 shows a slight decline in Q6 (Feedback), Q7 (Visual Appeal), and Q9 (Color Comfort). This can be attributed to some users either not noticing the haptic feedback or finding it unremarkable, as well as a few comments suggesting the design felt visually plain or lacking color. Q11 (Text Readability) shows a noticeable increase when Design One was experienced second, indicating that users in Group B found it easier to read, especially in contrast to the less readable text in Design Two.
Figure 9: Design One survey results.
In contrast, Design Two performed worse overall, with noticeable effects based on the order of exposure. When presented first, it received an average score of 3.4 across all questions. However, when experienced after Design One, the average rating of Group A dropped significantly to 2.9, indicating that prior exposure to a better-designed interface made users more aware of the flaws in the second design. As shown in Figure 3.1, the most significant drops were observed in Q2 (Comfort), Q3 (Ease of Interaction), and Q4 (Organization of Elements), suggesting that users found Design Two notably less comfortable, harder to use, and poorly structured when compared directly to a well-designed interface.
Figure 10: Design Two survey results.
Mean survey scores (1–5) by question showing that Design One stays consistently high regardless of whether users saw it first or second, whereas Design Two’s scores fall significantly, particularly for comfort (Q2), interaction ease (Q3), and element organization (Q4), when experienced after the superior design.
3.2 Overall Preference
@@ 6 Conclusion
Aknowledgements
This research was financially supported by the TUM Campus Heilbronn Incentive Fund 2024 of the Technical University of Munich, TUM Campus Heilbronn. We gratefully acknowledge their support, which provided the essential resources and opportunities to conduct this study.
[…]